OpenStudy (anonymous):

I'm having trouble with Snell's Law. Here is one of the problems A ray of light travels from lead to vacuum. If the ray of light in the vacuum makes an angle of 71.0 degrees to the normal, what is the angle of incidence in the lead?

OpenStudy (anonymous):

we have to use this equation\[n _{1}\sin \theta _{1}=n _{2}\sin \theta _{2}\]

OpenStudy (anonymous):

index of refraction of lead is required